determin the equasion of a line tha passes through (6,18) and (-12,3)
Answer by rfer(16322)

m=(3-18)/(-12-6)
m=-15/-18
m=5/6
---------------
y=mx+b
3=5/6*-12+b
3=-10+b
b=-13
---------------
y=5/6x-13